# MOTI: A Drug-Target Interaction Graph For Inductive Link Prediction

### v1

@@ Abstract
Drug-target interaction (DTI) prediction is crucial for identifying new therapeutics and detecting mechanisms of action. While structure-based methods accurately model physical interactions between a drug and its protein target, cell-based assays such as Cell Painting can better capture complex DTI interactions. This paper introduces MOTI, a Morphological cOmpound Target Interaction Graph dataset that comprises Cell Painting features for genes and compounds along with their relationships extracted from seven publicly available databases. We provide random, cold-source (new drugs), and cold-target (new genes) data splits to enable rigorous evaluation under realistic use cases. Our benchmark results show that graph neural networks that use Cell Painting features consistently outperform those that learn from graph structure alone, feature-based models, and topological heuristics. MOTI accelerates both graph ML research and drug discovery by promoting the development of more reliable DTI prediction models. MOTI resources are available at https://github.com/carpenter-singh-lab/motive.
@@ 1 Introduction
High quality graph benchmarking datasets propel graph machine learning (ML) research. Providing a diversity of domains, tasks, and evaluation methods, they allow for rigorous and extensive explorations of structured learning methods. Still, gaps remain in the space of scalability, network sparsity, and generalizability under realistic data splits [1]. These challenges are particularly relevant in the biological domain. The representation of the rich heterogeneity between entities—compounds, genes, proteins, diseases, phenotypes, side effects, and more—is a nontrivial task due to their varied units and terminology and highly complex relational structure. This makes biological data an apt, challenging, and bettering application for graph ML.
Next, effectively predicting drug-target interactions (DTIs), the relationships between chemical compounds and their protein targets, remains a pressing research area due to its relevance to drug discovery, drug repurposing, understanding side effects, and virtual screening. The DTI task is challenging due to the shortage of clean perturbational data and unspecificity of these interactions. Even as structure-based methods such as AlphaFold3 are increasingly accomplished at making DTI predictions, they are mainly based on molecular characteristics [2, 3]. Experimental data uniquely captures complex biological interactions; the morphological profiles, feature vectors which capture a cell’s appearance, from the Cell Painting (CP) assay have been shown to model mechanism of action, toxicity, and more [4, 5].
@@ 2 Related work
a trade off between scale and availability of node features. The Open Graph
Benchmark Library thus contributed ogbl-ppa, ogbl-collab and
ogbl-citation2, all large-scale, feature-based link prediction datasets.
The node features in each of these datasets are 58- or 128-dimensional and are a
one-hot representation of the protein type in ogbl-ppa or Word2Vec-based
representations of an author’s publications or of a paper’s contents in
ogbl-collab and ogbl-citation2 respectively. The benchmarking
@@
results showed a continued reliance on model learning from previous connections rather
@@
than features, as evidenced by the high performance of Node2Vec embeddings,
@@
and indicated a need for richer features. The authors also reported that the
graph neural network (GNN) models underperformed in the link prediction task
@@
when using mini-batch training rather than whole batch and called for
improvement in this area for future scalability when learning on large datasets.
In addition, the latter two datasets split their data by time metadata
@@
associated with each edge, and did not explore cold start splits. Evidently, the field still requires graph datasets that 1) are large scale, 2)
include information rich features, 3) are accompanied by graph-based data splits
@@
(not metadata-based), and 4) are flexibly trained with mini-batch sampling. We
@@
prioritized all four of these goals during the curation of MOTI and in our
experimental design.
@@
As Knowledge Graphs (KGs) have emerged as powerful tools for representing and learning from network-based data[10], they have been often applied to biological and chemical tasks such as drug discovery, structure prediction, and DTI [11, 12, 13]. Recent surveys on graph-based methods for DTI prediction [14, 15] show few efforts represent drugs and genes with external features. Two related approaches are Schwarz et al. [16], which used gene expression as input to the last layer in a late-fusion manner and Balamuralidhar et al. [17], which used topological features as node representations. While Balamuralidhar et al. [17] does enable inductive predictions on newly connected nodes, it fails to make inferences on completely isolated nodes in cold start scenarios. To our knowledge, no graph-based dataset exists where the features of compounds and gene nodes are represented by their morphological profiles.
@@
Meanwhile, in biological image analysis, fluorescence labeling of cells now allows for the visualization of cell morphology, internal structures, and processes at unprecedented spatial and temporal resolution. Imaging precisely captures the changes to a cell after it has undergone a chemical or genetic perturbation and sheds light on relationships such as DTI, functions, and mechanisms. Additionally, the recent publication of the JUMP Cell Painting dataset [6] now provides chemical and genetic perturbational profiles for the DTI task. In recent applications of CP to DTI, Rohban et al. [18] linked drugs and targets by correlation based matching of their morphological features, which may not capture higher-order relationships. Herman et al. [4] developed a deep learning model to predict assay activity based on the chemical structure and morphological features. Although they included graph-based methods to represent molecular structures, they did not exploit the network connectivity across compounds and genes. Recently, Iyer et al. [19] formulated the DTI task as a binary classification of gene-compound pairs under different data splits and developed a transformer-based learning approach to predict drug targets from Cell Painting profiles. Although this approach is similar to our proposed setup, their dataset contains a reduced number of nodes, 302 compounds and 160 genes, and does not include gene-gene and compound-compound interactions.
@@ 3.1 Morphological profiles extraction
Every compound and gene (each of which produces a particular protein) is represented with its image-based profile from the JUMP CP dataset [6]. JUMP CP images capture a cell’s morphology after being perturbed by a chemical compound or a genetic edit. Each type of genetic perturbation, CRISPR knockouts or ORF over-expressions, involves a different set of genes. We created and published a version of MOTI for each gene set, but chose to use the ORF genes for the analysis presented in this paper due to slightly stronger downstream performance. The morphological profiles were extracted from each image in the JUMP CP dataset using CellProfiler [20], a software that segments individual cells in the image and measures thousands of features for each cell. All of the profiles associated with a perturbation are aggregated, batch corrected, and preprocessed according to the protocols laid out in Arevalo et al. [21] and Chandrasekaran et al. [6]. After the correction and preprocessing steps, each compound and gene is represented by a 737-dimensional and 722-dimensional vector respectively.
@@ 3.2 Annotation collection
We aggregated the compound-compound, compound-gene and gene-gene annotations in MOTI from seven publicly available databases (Table 1). Our choice followed the comprehensive review of Bonner et al. [27], which categorizes relevant biomedical datasets and KGs based on the entities they contain and how well they fit to specific tasks. We unified the compound IDs by mapping them to their core molecular skeleton identifiers (InChIKey) using the MyChem [28] and UniChem [29] databases and unified the genes IDs using their NCBI gene symbols 111https://www.ncbi.nlm.nih.gov/gene/. We then combined all of the pairwise interactions we collected across databases with the JUMP CP features, keeping only the pairs where both entities had existing features and using only the features for which we had interactions.
@@ 3.4 Data Splitting
We developed two different heuristics to split our graph into training, validation, and test sets based on the random split and cold start split originally defined for recommendation systems [7]. In the random split scenario, we used a 70/10/20 ratio to randomly select and split every edge into train, validation, and test sets and included the full set of and edges in training (Figure 1). In this case, the link prediction task is transductive, as the nodes are fixed from the start of training and the model will predict edges on entities it has learned on.
The cold start split allows for inductive link prediction, as it involves predicting edges where at least one node was not present during training. We applied the cold start split to either the source or target nodes, denoted by cold-source split or cold-target split. In cold-source split (second row of Figure 1), every source node in our graph is randomly labeled as either train, validation, or test in a 70/10/20 ratio. All edges are subsequently labeled in accordance with the label of its source node. Next, all edges are labeled by their most conservative labels: any edge with at least one test source is labeled test, any edge with no test sources but at least one validation source is labeled validation, and any edge with two train sources is labeled train. All edges are included in training. This data split prevents data leakage of any kind and enforces that at validation or test time, the model has never seen any source node in the edges. Additionally, negative edges are sampled for each batch of data by our custom sampler class, built for inductive learning, which only considers the relevant left out source nodes and all of the unique target nodes in the edges. Details about our negative sampling algorithm can be found in Section 3.5.
@@ 3.5 Negative Sampling Algorithm
We used a negative sampling ratio () of 1:1 during training and 1:10 during testing. For random split, our algorithm samples the negative edges between all unique source and target nodes within the edges in the batch. For cold-source split, the head of the negative edge is sampled from the unique sources in the supervision edges in the batch, and the tail of the negative edge is sampled from the unique targets in the message-passing or supervision edges in the batch. This method ensures the distributions of source and target nodes are identical for positive and negative supervision edges. Algorithm 1 details the cold-source split negative sampling procedure. The process is symmetrical for leave-out-target.
@@ 3.6 Models
GraphSAGE (Graph SAmple and aggreGatE) [5] is a GNN for inductive representation learning on large graphs. Our proposed model, GraphSAGECP, adapts GraphSAGE to use the JUMP Cell Painting features as the nodal representations. First, the input embeddings of every source and target node are initialized and fixed as their respective feature vectors and . Two linear layers then transform the source and target node embeddings into the same embedding space. Here, the embeddings pass through two SAGEConv layers, separated by a leaky ReLU activation function for nonlinearity. Each SAGEConv layer contains two weight matrices, which respectively weigh the feature vector of the node of interest and the average of the feature vectors from the neighbors of the node of interest. For isolated nodes, there is no signal from the neighbors. Essentially, GraphSAGECP convolves the feature vectors of the neighbor nodes within a radius of 2 to compute the hidden embedding for each node. An additional skip connection feeds the output of the first GraphSAGE layer (an aggregation of the feature vectors in the neighborhood of radius 1) into the final embedding, prioritizing shorter distance neighbors. Finally, a classifier head outputs the sigmoid-transformed dot product of the embeddings for each source and target node pair in the supervision edge set. Algorithm 2 in Appendix A further describes the forward pass of our model to make the link predictions.
@@
We benchmarked our model with a featureless graph-based model that randomly initializes the source and target input embeddings. Other than the initialization and weight updates for the embedding layers, this model, GraphSAGEembs, shares the same architecture as GraphSAGECP. Importantly, GraphSAGEembs does not have informative feature vectors for each source and target node at the start of training. We also applied one baseline heuristic for topology-based link predictions, and two baseline models for feature-based link predictions. First, we used the shortest path between source and target nodes as proxy for their similarity to evaluate the predictive power of graph structure alone. Next, we implemented a Bilinear model and a Multilayer Perceptron (MLP), which both rely solely on node features to predict links. The Bilinear model learns a mapping to transform two sets of feature vectors into the same embedding space, then computes the similarity between the vectors. The MLP learns hidden embeddings of the feature vectors via two fully connected layers and ReLU activation functions for nonlinearity, then uses a Bilinear head to make link predictions. These latter models ignore the known relational information between entities when predicting links. Across the five models, we could readily evaluate the performance enhancements coming from adding node features or graph structure and determine the most useful learning approach for each of our scenarios. See Appendix B for experimental details and evaluation metric choices.
@@ 4.1 DTI prediction improves with CP features
| Model | Input | F1 | Hits@500 | Precision@500 |
| GraphSAGECP | Graph+CP | 0.5947 0.008 | 0.3955 0.008 | 0.9716 0.005 |
| GraphSAGEembs | Graph | 0.5345 0.004 | 0.3165 0.006 | 0.9456 0.008 |
| MLP | CP | 0.2633 0.023 | 0.05230.002 | 0.3552 0.024 |
| Bilinear | CP | 0.1779 0.004 | 0.0281 0.001 | 0.2260 0.009 |
| Shortest Path | Graph | — | 0.0025 | 0.0 |
First, we evaluated all of our models on transductive link predictions (random split). We used the shortest path heuristic, a non-learning baseline, to predict links based on graph topology (last row of Table 3). We scored each positive and negative edge in our test set by the shortest path length between the two nodes and computed rank based metrics based on these scores. This heuristic fails to effectively predict links on MOTI, proving the task nontrivial and indicating a need for learning-based methods. The low prediction scores from the Bilinear and MLP models (rows 3 and 4 of Table 3) also indicate that predicting edges purely based on the similarity of source and target node features is inadequate. Still, these models significantly outperform the non-learning baseline, signaling the information contained in the node features.
Next, we see a large score increase across all metrics between the MLP and GraphSAGEembs model. The predictions made by the learned representations of source and target nodes from topology alone achieves a Precision@500 score of 94.56%, which highlights the richness of the relational information between nodes. Finally, we see the highest Precision@500 of 97.16% from GraphSAGECP, demonstrating that adding CP node features to graph structure benefits transductive link prediction. However, even the strongest model, GraphSAGECP, only achieves a 59.47% F1 score and a 39.55% Hits@500 score, indicating opportunity for improvement.

### v2

@@ Abstract
Drug-target interaction (DTI) prediction is crucial for identifying new therapeutics and detecting mechanisms of action. While structure-based methods accurately model physical interactions between a drug and its protein target, cell-based assays such as Cell Painting can better capture complex DTI interactions. This paper introduces MOTI, a Morphological cOmpound Target Interaction Graph dataset comprising Cell Painting features for genes and compounds, along with their relationships extracted from seven publicly available databases. We provide random, cold-source (new drugs), and cold-target (new genes) data splits to enable rigorous evaluation under realistic use cases. Our benchmark results show that graph neural networks that use Cell Painting features consistently outperform those that learn from graph structure alone, feature-based models, and topological heuristics. MOTI accelerates both graph ML research and drug discovery by promoting the development of more reliable DTI prediction models. MOTI resources are available at https://github.com/carpenter-singh-lab/motive.
@@ 1 Introduction
High-quality graph benchmarking datasets propel graph machine learning (ML) research. Providing a diversity of domains, tasks, and evaluation methods, they allow for rigorous and extensive explorations of structured learning methods. Still, gaps remain in the areas of scalability, network sparsity, and generalizability under realistic data splits [1]. These challenges are particularly relevant in the biological domain. The representation of the rich heterogeneity between entities—compounds, genes, proteins, diseases, phenotypes, side effects, and more—is a nontrivial task due to their varied units and terminology and highly complex relational structure. This makes biological data an apt, challenging, and bettering application for graph ML.
Next, effectively predicting drug-target interactions (DTIs), the relationships between chemical compounds and their protein targets, remains a pressing research area due to its relevance to drug discovery, drug repurposing, understanding side effects, and virtual screening. The DTI task is challenging due to the shortage of clean perturbational data and nonspecificity of these interactions. Even as structure-based methods such as AlphaFold3 are increasingly accomplished at making DTI predictions, they are mainly based on molecular characteristics [2, 3]. Experimental data uniquely captures complex biological interactions; the morphological profiles, feature vectors that capture a cell’s appearance, from the Cell Painting (CP) assay have been shown to model the mechanism of action, toxicity, and additional properties of compounds [4, 5].
@@ 2 Related work
a trade-off between scale and availability of node features. The Open Graph
Benchmark Library (OGBL) thus contributed ogbl-ppa
, ogbl-collab
and ogbl-citation2
, all large-scale, feature-based link prediction
datasets. The node features in each of these datasets are 58- or 128-dimensional
and are a one-hot representation of the protein type in ogbl-ppa
or
Word2Vec
-based representations of an author’s publications or of a
paper’s contents in ogbl-collab
and ogbl-citation2
respectively.
@@
The benchmarking results showed a continued reliance on model learning from
@@
previous connections rather than features, as evidenced by the high performance
of Node2Vec
embeddings in OGBL tasks, and indicated a need for richer
@@
features. The authors also reported that the graph neural network (GNN) models
@@
underperformed in the link prediction task when using mini-batch training rather
than whole batch and called for improvement in this area for future scalability
when learning on large datasets. In addition, the latter two datasets split
their data by time metadata associated with each edge, and did not explore cold
@@
start splits. Evidently, the field still requires graph datasets that 1) are
large-scale, 2) include information-rich features, 3) are accompanied by
@@
graph-based data splits (not metadata-based), and 4) are flexibly trained with
@@
mini-batch sampling. We prioritized all four goals during the curation of
MOTI and in our experimental design.
@@
As Knowledge Graphs (KGs) have emerged as powerful tools for representing and learning from network-based data [10], they have often been applied to biological and chemical tasks such as drug discovery, structure prediction, and DTI [11, 12, 13]. Recent surveys on graph-based methods for DTI prediction [14, 15] show few efforts represent drugs and genes with external features. Two related approaches are Schwarz et al. [16], which used gene expression as input to the last layer in a late-fusion manner, and Balamuralidhar et al. [17], which used topological features as node representations. While Balamuralidhar et al. [17] does enable inductive predictions on newly connected nodes, it fails to make inferences on completely isolated nodes in cold start scenarios. To our knowledge, no graph-based dataset exists where the features of compounds and gene nodes are represented by their morphological profiles.
@@
Meanwhile, in biological image analysis, fluorescence labeling of cells now allows for the visualization of cell morphology, internal structures, and processes at unprecedented spatial and temporal resolution. Imaging precisely captures the changes to a cell after it has undergone a chemical or genetic perturbation and sheds light on relationships such as DTI, functions, and mechanisms. Additionally, the recent publication of the JUMP Cell Painting dataset [6] now provides chemical and genetic perturbational profiles for the DTI task. In recent applications of CP to DTI, Rohban et al. [18] matched compounds to a small set of genes based on morphological feature vector similarities, and we will extend this work by using machine learning to capture non-linear compound-gene relationships for a much larger gene set. Next, Herman et al. [4] developed a deep learning model to predict toxicity assays from chemical structures and morphological profiles but did not exploit the connectivity network of compounds and gene interactions. Our method builds on this approach by incorporating the morphological profiles of both compounds and genes in a graph and using the message-passing framework to leverage such network connectivity. Recently, Iyer et al. [19] formulated the DTI task as a binary classification of gene-compound pairs under different data splits and developed a transformer-based learning approach to predict drug targets from Cell Painting profiles. Although this approach is similar to our proposed setup, their dataset contains fewer nodes, 302 compounds and 160 genes, and does not include gene-gene and compound-compound interactions.
@@ 3.1 Morphological profiles extraction
Every gene (each of which produces a particular protein) and compound is represented with its image-based profile from the JUMP CP dataset [6]. JUMP CP images capture a cell’s morphology after being perturbed by a chemical compound or a genetic edit. Each type of genetic perturbation, CRISPR knockouts or ORF over-expressions, involves a different set of genes. We created and published a version of MOTI for each gene set but chose to use the ORF genes for the analysis presented in this paper due to slightly stronger downstream performance. The morphological profiles were extracted from each image in the JUMP CP dataset using CellProfiler [20], a software that segments individual cells in the image and measures thousands of features for each cell. Then, the data was prepared according to the protocols in Arevalo et al. [21] and Chandrasekaran et al. [6], which extensively optimize separate pipelines for compound and gene perturbations.
For compound perturbations, we first filtered out features with low variance, then applied median absolute deviation normalization, a rank-inverse normal transformation (INT), and Harmony [22] to reduce the batch effects. We then selected final features based on correlation analysis. For genetic perturbations, we subtracted the mean vector per well from each feature vector to account for well position effects, replaced the INT transformation with an outlier removal step, and aggregated the replicates (usually five) of each perturbation using median profiles to make the representations robust to low-quality images. In our experience, these comprised less than of the data and were uncorrelated across replicates on different plates. After the correction and preprocessing steps, each compound and gene is represented by a 737-dimensional and 722-dimensional vector, respectively. The processing pipelines for the morphological profiles are available at https://github.com/broadinstitute/jump-profiling-recipe/tree/v0.1.0.
@@ 3.2 Annotation collection
We aggregated the compound-compound, compound-gene and gene-gene annotations in MOTI from seven publicly available databases listed in Table 1. Our choice followed the comprehensive review of Bonner et al. [28], which categorizes relevant biomedical datasets and KGs based on the entities they contain and how well they fit to specific tasks. We unified the compound IDs by mapping them to their core molecular skeleton identifiers (InChIKey) using the MyChem [29] and UniChem [30] databases and unified the gene IDs using their NCBI gene symbols 111https://www.ncbi.nlm.nih.gov/gene/. We then combined all of the pairwise interactions we collected across databases with the JUMP CP features, keeping only the pairs where both entities had existing features and using only the pairs for which we had interactions.
@@ 3.4 Data splitting
We developed two different heuristics to split our graph into training, validation, and test sets based on the random split and cold start split originally defined for recommendation systems [7].
In the random split scenario, we used a 70/10/20 ratio to randomly select and split every edge into train, validation, and test sets and included the full set of and edges in training (Figure 1). In this case, the link prediction task is transductive, as the nodes are fixed from the start of training and the model will predict edges on entities it has learned on. Negative edges are sampled for each batch of data by randomly selecting source-target pairs that are not in the edges.
The cold start split allows for inductive link prediction, as it involves predicting edges where at least one node was not present during training. We applied the cold start split to either the source or target nodes, denoted by cold-source split or cold-target split. In cold-source split (second row of Figure 1), every source node in our graph is randomly labeled as either train, validation, or test in a 70/10/20 ratio. All edges are subsequently labeled in accordance with the label of its source node. Next, all edges are labeled by their most conservative labels: any edge with at least one test source is labeled test, any edge with no test sources but at least one validation source is labeled validation, and any edge with two train sources is labeled train. All edges are included in training. This data split prevents data leakage of any kind and enforces that at validation or test time, the model has never seen any source node in the edges. Additionally, negative edges are sampled for each batch of data by our custom sampler class, built for inductive learning, which only considers the sampled source nodes and all of target nodes, in the edges. Details about our negative sampling algorithm can be found in Section 3.5.
@@ 3.5 Negative sampling algorithm
We used a negative sampling ratio () of 1:1 during training and 1:10 during testing. For random split, our algorithm samples the negative edges between all unique source and target nodes within the edges in the batch. For cold-source split, the head of the negative edge is sampled from the unique sources in the supervision edges in the batch, and the tail of the negative edge is sampled from edges in the batch. Algorithm 1 details the cold-source split negative sampling procedure. The process is symmetrical for leave-out-target.
@@ 3.6 Models
We experimented with three different types of GNN convolutional layers in our model, each with a unique learning algorithm. The algorithms differ in how they incorporate the features of neighbor nodes into a node’s learned representation. GraphSAGE (Graph SAmple and aggreGatE) [5] achieves inductive representation learning on large graphs by sampling from neighbor node representations and applying an aggregation function (e.g. a weighted average) to compute each node’s hidden embedding. GIN (Graph Isomorphism Network) [32] adds a Multilayer Perceptron (MLP) to represent the composition of neighbor node features and achieves discriminative power equal to the Weisfeiler Lehman graph isomorphism test. Finally, GATv2 (Graph Attention Network v2) [33] implements dynamic graph attention which weights neighbor representations according to each query of interest. We constructed separate models using each of these GNN algorithms (GraphSAGECP, GINCP, and GATv2CP) as the convolution layer, and we used the JUMP Cell Painting features as the nodal representations for each. We chose to only initialize the node features using their image-based features (as opposed to chemical structure-based or protein structure-based) so that compounds and genes would share a modality of representation. This ensures similar statistical properties between node types and makes it easier to discover cross-modal relationships [34].
@@
All three GNN models share the same architecture. First, the input embeddings of every source and target node are initialized and fixed as their respective feature vectors and . Two linear layers then transform the source and target node embeddings into the same embedding space. These transformed embeddings pass through two GNN layers (chosen from one of the three graph convolutional algorithms above), separated by a leaky ReLU activation function for nonlinearity. Broadly, each GNN layer combines the feature vector of the node of interest with some aggregation of the feature vectors from the neighbors of the node of interest. The aggregation function changes according to the three GNN algorithms. Isolated nodes rely solely on CP features due to the lack of neighboring signals. The two GNN layers indicate that the feature vectors of neighbor nodes within a radius of 2 will be used to compute the hidden embedding for each node. An additional skip connection feeds the output of the first GNN layer (an aggregation of the feature vectors in the neighborhood of radius 1) into the final embedding, prioritizing shorter distance neighbors. Finally, a classifier head outputs the sigmoid-transformed dot product of the embeddings for each source and target node pair in the supervision edge set. Algorithm 2 in Appendix A further describes the forward pass of our GNNs to make the link predictions.
@@
We benchmarked our model with a featureless graph-based model that randomly initializes the source and target input embeddings. We opted to use the simple GraphSAGE convolutional layer in our model because all of the node features are random vectors in this case. This model, GraphSAGEembs, shares the same architecture as the other GNNs. Importantly, GraphSAGEembs does not have informative feature vectors for each source and target node at the start of training. We also applied one baseline heuristic for topology-based link predictions, and two baseline models for feature-based link predictions. For the former, we used the shortest path between source and target nodes as a proxy for their similarity to evaluate the predictive power of graph structure alone. For the latter, we implemented a Bilinear model and an MLP, which both rely solely on node features to predict links. The Bilinear model learns a mapping to transform two sets of feature vectors into the same embedding space, then computes the similarity between the vectors. The MLP learns hidden embeddings of the feature vectors via two fully connected layers and ReLU activation functions for nonlinearity, then uses a Bilinear head to make link predictions. These latter models ignore the known relational information between entities when predicting links. Across the seven models, we could readily evaluate the performance enhancements coming from adding node features or graph structure and determine the most useful learning approach for each of our scenarios. See Appendix B for experimental details and evaluation metric choices.
@@ 4.1 DTI prediction improves with CP features
Input | Model | F1 | Hits@500 | Precision@500 |
GINCP | 0.5238 0.040 | 0.4552 0.017 | 0.9920 0.005 | |
Graph+CP | GATv2CP | 0.4169 0.007 | 0.2219 0.018 | 0.8452 0.045 |
GraphSAGECP | 0.3836 0.029 | 0.2637 0.023 | 0.9056 0.012 | |
CP | MLP | 0.3829 0.008 | 0.2545 0.011 | 0.8456 0.015 |
Bilinear | 0.1703 0.003 | 0.0213 0.001 | 0.1812 0.008 | |
Graph | GraphSAGEembs | 0.34560.006 | 0.22540.013 | 0.84760.029 |
Shortest Path | — | 0.0025 | 0.0 |
First, we evaluated all of our models on transductive link predictions (random split). We used the shortest path heuristic, a non-learning baseline, to predict links based on graph topology (last row of Table 3). We scored each positive and negative edge in our test set by the shortest path length between the two nodes and computed rank based metrics based on these scores. This heuristic fails to effectively predict links on MOTI, proving the task nontrivial and indicating a need for learning-based methods. The low prediction scores from the Bilinear and MLP models (rows 4 and 5 of Table 3) also indicate that predicting edges purely based on the similarity of source and target node features is inadequate. Still, these models significantly outperform the non-learning baseline, signaling the information contained in the node features.
Next, we see a large score increase from the feature-based benchmarks to the GNN models, which make use of the graph structure and CP node features. The predictions made by the learned representations of source and target nodes from topology alone (GraphSAGEembs) achieve a Precision@500 score of 84.76%, which highlights the richness of the relational information between nodes. Finally, the GINCP model achieves the best performance across metrics, obtaining a Hits@500 score of 45.54% and a Precision@500 score of 99.20%. This result supports our hypothesis and demonstrates that adding CP node features to graph structure benefits transductive link prediction.